Two large containers A and B of the same size are filled with different fluids. The fluids in containers A and B are maintained at 0° C and 100° C, respectively. A small metal bar, whose initial temperature is 100° C, is lowered into container A. After 1 minute the temperature of the bar is 90° C. After 2 minutes (since being lowered into container A) the bar is removed and instantly transferred into the other container. After 1 minute in container B the temperature of the bar rises 10°. How long, measured from the start of the entire process, will it take the bar to reach 99.5° C? (Round your answer to two decimal places. Assume the final temperature being asked for is reached while the bar is container B.)

t = ______ min
Answer #1

Ô solution : Z Given that Container A has a Temperature of OC and containeo B hao a Temperatute of 100°C Foomula : - nedtonsthat al we know The ambient temperature .foo this case is Tm=0. The Tempecabute of contineo A, finally t: I, The baris Tempeen[.993)-tln() ki In10.995) en 1966 t: (-50125#100) -0-105360515 % =0.047375 minutes The bat has a temperature of 99.5% neat,6:12 โ: 91 solve Foo K 91-100=-elok +9=41acht 91:-190 xBf 100 olo cekti) en la alk 56. The Equalion modeling the temperature

I HOPE THIS ANSWER WILL BE HELPFUL TO YOU

PLEASE LIKE THIS ANSWER

THANK YOU.